TIP: Click on subject to list as thread! ANSI
echo: sb-nasa_news
to: All
from: Hugh S. Gregory
date: 2003-02-16 23:14:00
subject: 1\31 Pt-1 ISS On Orbit Status 31-01-2003

This Echo is READ ONLY !   NO Un-Authorized Messages Please!
 ~~~~~~~~~~~~~~~~~~~~~~~~   ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
1\31 ISS On-Orbit Status 31-01-2003
Part 1 of 2

ISS On-Orbit Status 1/31/03

All ISS systems continue to function nominally, except as noted
previously or below.

Wake-up was at the regular 1:00am EST, followed by post-sleep 
activities and breakfast. Begin of sleep time: 4:30pm.

FE-1 Nikolai Budarin and CDR Kenneth Bowersox continued working on the
Progress 9P spacecraft in preparation for its undocking tomorrow 
morning.  After power-down of the cargo vehicle's US-21 matching unit 
by MCC-M/TsUP and disconnection of the BITS 2-12 onboard telemetry 
measurement system from it by Budarin, he and Bowersox removed the 
no-longer-needed US-21 in its container box. [The matching unit and 
associated commutator gear provide the electronic interface between 
the Service Module (SM) and Progress for SM computer control of 
Progress propulsion, i.e., during ISS reboost maneuvers.  The avionics 
will be recycled on a future cargo flight. Progress undocking and 
post-separation control will be handled by the drone's pre-programmed 
automated sequencing system.]

After the R&R, Budarin and Bowersox conducted final undocking 
preparations.  Separation command will nominally be given tomorrow 
morning at 10:59am EST, followed by deorbit burn ignition at 2:10pm 
and atmospheric entry beginning at 2:47pm. Plume observations will be 
conducted from the SM with the infrared spectrometer of the Russian 
geophysical GFI-1 "Relaksatsiya" experiment. [Nikolai activated 
Progress and removed the air duct between it and the SM transfer 
tunnel (PrK) at Zvezda's aft end. QD (quick disconnect) screw clamps 
were removed and the Progress/SM interface was video-surveyed with the 
U.S. Sony PD-100 DVCAM for subsequent downlink to and inspection by 
TsUP. Hatch closure proceeded at about 9:30am EST (which also served 
as decision point for final irreversible launch preparations at 
Baikonur for Progress 10P), followed by an interface leak check of the 
Progress/PrK vestibule, during which Russian thrusters were disabled.]

Launch countdown for Progress 10P (M-247) at Baikonur continues toward
liftoff on 2/2 (Sunday), at 7:59am EST. Docking is set for 2/4 
(Tuesday), at 9:48am. [Since the ISS solar arrays will be feathered 
during the 2-hr.  docking window (i.e., not tracking the Sun), the 
expected power generation decrease will require some temporary 
powerdowns in the U.S. segment (USOS) and of one USOS-to-RS (Russian 
segment) SNT-50MP voltage and current stabilizer unit (transformer). 
Details are still in work.]

As POC (Payload Operations Center) continues to troubleshoot the 
recent failure of the ER3 ARIS (EXPRESS rack 3/active rack isolation 
system) calibration, Bowersox continued ER work, today loading the ER2 
laptop (ELC) with the new software, but not yet the RIC (rack 
interface controller). [The new load was expected to help troubleshoot 
why the laptop is not communicating with the rack. If it is due to a 
corrupt file, as suspected, the s/w load will fix the problem. If it 
is a hardware problem, however, the load may not complete 
successfully. Prior to the load, Don Pettit needed to swap out the 
ELC's hard drive with a spare (with Windows NT, instead of Windows 95 
or 98).]

Today, the highly successful EarthKAM operation came to an end for 
Increment 6. FE-2/SO Pettit deactivated the camera system and stowed 
it, reconfiguring the cabling to get the Lab window ready for the 
installation of the WORF (window observational research facility), a 
modified ISPR (international standard payload rack) due to arrive in 
early March on STS-114/ULF-1.

Budarin checked up on the MO-21 "Ecosfera" air sampler and incubation
equipment, monitoring colony growth for atmospheric microbial 
analysis.  [MO-21 determines microbial contamination of the ISS 
atmosphere, specifically the total bacterial and fungal microflora 
counts and microflora composition according to morphologic criteria of 
microorganism colonies.]

Later, he was scheduled to tag up with an Uragan ("hurricane") 
specialist to discuss the new earth imaging sessions planned for 
tomorrow and Sunday.

After completion of the regeneration process on the second Metox 
(metal oxide) canister, which can now be used again in the next EVA 
operations as a carbon dioxide (CO2) filter, the crew monitored the 
deactivation of the U.S.  CDRA (CO2 removal assembly). The ground also 
reset the Airlock temperature setpoint back up to 70F (21 degC).

In the RS, the Vozdukh CO2 scrubber remains powered off for a 
long-planned efficiency test of the Russian LiOH (lithium hydroxide) 
cartridges.  [Intended to support extension of certified life time of 
the backup CO2 absorbers, the activation of one cartridge was to test 
its capability to remove CO2 from the cabin without Vozdukh or CDRA 
running.]

Recent ppCO2 (CO2 partial pressure) values have consistently been 
higher in the USOS than in the SM, as indicated by the Russian sensor 
system. For the LiOH canister efficiency test in the RS, it was 
suggested to Moscow to use the portable U.S. CDMK (CO2 monitoring kit) 
instrument instead, to ensure that Lab ppCO2 concentration does not go 
over the level of 6.0 mmHg.

CDR Bowersox set up the video equipment for recording the crew's 
workout on the RED (resistive exercise device). All three crewmembers 
were scheduled to record their anaerobic workout, and the video camera 
was to be stowed after the last session of the day. Later, Don Pettit 
performed the regular weekly downloading of exercise files from the 
TVIS treadmill, RED and CEVIS bike to the MEC (medical equipment 
computer).

Nikolai Budarin completed his daily checkup of the Russian BIO-5
Rasteniya-2/Lada-2 ("Plants-2") plant growth experiment.

Don Pettit performed the regular daily tasks of SOSh life support 
systems maintenance and IMS database delta file preparation, while Ken 
Bowersox did the Lab payload status checkup (PCG-STES010).

Sox also began the first part of an extensive "pantry audit" of 
on-board provisions, entailing such item categories as office 
supplies, towels and napkins, printer supplies, batteries, sanitary 
hygiene provisions, and hygiene skip-cycle supply (in case of delayed 
resupply). Due to Increment 7/ ULF-1 packing deadlines, today's audit 
gave priority to the office supplies part. [The audit is intended to 
verify the complement of U.S. crew provisions onboard, to allow MCC-H 
to determine necessary resupply items and which items are not being 
used and are not needed, and to support Increment 7/ULF-1 packing as 
well as Increment 8 manifesting.]

Pending final signature of the jointly prepared water venting protocol 
by the Russian side, the next Lab condensate dump will take place next 
Friday (2/7), performed non-propulsively (i.e., from two vents) in the 
current LVLH attitude, as opposed to the more regularly used XPOP 
attitude. If the agreement is not signed in time, the excess 
condensate will be dumped into a CWC (contingency water container).

Troubleshooting of the tripped RPCM (remote power controller module)
continues. If the VTR1 (video tape recorder #1), which it powers, is 
found not at fault, the RPCM will be changed out. If the VTR is the 
cause of the trip, there are two VTR spares on the ground, ready to go 
if needed to replace it (typically on 12A, but this may change).

(continued)

---
* Origin: SpaceBase[tm] Vancouver Canada [3 Lines] 604-473-9357 (1:153/719)
SEEN-BY: 633/267 270
@PATH: 153/719 715 7715 140/1 106/2000 633/267

SOURCE: echomail via fidonet.ozzmosis.com

Email questions or comments to sysop@ipingthereforeiam.com
All parts of this website painstakingly hand-crafted in the U.S.A.!
IPTIA BBS/MUD/Terminal/Game Server List, © 2025 IPTIA Consulting™.